# SaVe-Tag: Semantic-aware Vicinal Risk Minimization for Long-Tailed Text-Attributed Graphs

### v3

@@ 5.3 Necessity of Confidence-based Edge Assignment (RQ2)
Figure 5: Top-k edge selection.
We instantiate Definition 4.7 using a top–$k$ rule, where each synthetic node retains its $k$ highest-confidence edges. The value of $k$ is set to #$\#augmented\_nodes\times n$, and we vary $n$ to study its effect on node isolation and classification performance. Figure 5a shows that the number of isolated nodes ($\log$-scaled) drop as $\log(n)$ increases. Figure 5b reveals a bell-shaped accuracy curve with respect to $\log(n)$, measured relative to $n{=}1$. These trends support our hypothesis:
  • Too small $k$rightarrow isolation. Many synthetic nodes become singletons (Theorem 4.9), so the minority manifold is under-populated and performance suffers.
  • Too large $k$rightarrow over-smoothing. Excessive connectivity leads to repeated averaging in Equation (1) and pushing node representations towards a global mean, resulting in over-smoothing [34].
@@
We compare two edge assignment strategies for augmented nodes: edge duplication from the target node (without the ‘${C}$’ subscript) and edge assignment with our confidence-based function (denoted by ‘${C}$’). Table 2 shows that our confidence-based edge assignment consistently outperforms naive edge duplication under LLM-based interpolation. A similar trend holds for numeric interpolation, with the exception of Cora and Photo, likely due to the generation of off-manifold samples that our confidence mechanism successfully isolates.
Dataset Method SaVe-TagS SaVe-TagM SaVe-TagO F1 $\uparrow$ (%) ACC $\uparrow$ (%) F1 $\uparrow$ (%) ACC $\uparrow$ (%) F1 $\uparrow$ (%) ACC $\uparrow$ (%) Cora Origin 70.40 + 0.0 72.46 + 0.0 70.40 + 0.0 72.46 + 0.0 70.40 + 0.0 72.46 + 0.0 Num 70.59 + 0.1 73.11 + 0.9 73.21 + 3.9 75.32 + 3.8 71.04 + 0.9 73.80 + 1.9 NumC 70.09 - 0.4 72.44 - 0.9 71.86 + 2.0 71.18 - 1.8 70.89 + 0.7 72.76 + 0.4 LLM 72.54 + 3.0 74.51 + 2.8 72.30 + 2.7 74.21 + 2.4 70.09 - 0.4 72.86 + 0.6 LLMC 74.05 + 5.2 76.36 + 5.4 70.54 + 0.2 72.51 + 0.1 73.21 + 4.0 75.32 + 4.0 PubMed Origin 47.79 + 0.0 55.09 + 0.0 47.79 + 0.0 55.09 + 0.0 47.79 + 0.0 55.09 + 0.0 Num 69.06 + 44.5 69.43 + 26.0 70.48 + 47.5 70.73 + 28.4 69.44 + 45.3 69.77 + 26.6 NumC 69.19 + 44.8 69.63 + 26.4 72.66 + 52.0 72.46 + 31.5 69.01 + 44.4 69.43 + 26.0 LLM 72.54 + 51.8 72.50 + 31.6 71.34 + 49.3 71.38 + 29.6 71.44 + 49.5 71.51 + 29.8 LLMC 76.43 + 59.6 76.35 + 38.6 76.18 + 59.4 76.26 + 38.4 72.96 + 52.7 72.80 + 32.1 Citeseer Origin 49.63 + 0.0 53.71 + 0.0 49.63 + 0.0 53.71 + 0.0 49.63 + 0.0 53.71 + 0.0 Num 62.69 + 26.3 65.89 + 22.7 63.52 + 28.0 67.24 + 25.2 63.12 + 27.2 66.28 + 23.4 NumC 65.19 + 31.4 68.79 + 28.1 64.10 + 29.2 68.89 + 28.3 64.75 + 30.5 68.24 + 27.1 LLM 66.51 + 34.0 69.19 + 28.8 66.93 + 34.9 69.69 + 29.8 66.99 + 35.0 69.58 + 29.5 LLMC 66.85 + 34.7 69.89 + 30.1 67.96 + 36.9 71.60 + 33.3 66.82 + 34.6 69.90 + 30.1 Photo Origin 54.99 + 0.0 51.69 + 0.0 54.99 + 0.0 51.69 + 0.0 54.99 + 0.0 51.69 + 0.0 Num 57.29 + 4.2 55.74 + 7.8 57.92 + 5.3 56.04 + 8.4 57.08 + 3.8 55.08 + 6.6 NumC 56.05 + 1.9 54.97 + 6.3 56.74 + 3.2 54.57 + 5.6 54.44 - 1.0 51.31 - 0.7 LLM 58.40 + 6.2 56.75 + 9.8 59.53 + 8.3 58.06 + 12.3 59.32 + 7.9 58.19 + 12.6 LLMC 59.20 + 7.7 58.65 + 13.5 59.45 + 8.1 58.60 + 13.4 63.62 + 15.7 66.17 + 28.0 Computer Origin 44.93 + 0.0 49.34 + 0.0 44.93 + 0.0 49.34 + 0.0 44.93 + 0.0 49.34 + 0.0 Num 48.90 + 8.8 54.47 + 10.4 49.47 + 10.1 53.85 + 9.1 49.47 + 10.1 55.78 + 13.0 NumC 49.31 + 9.7 54.69 + 10.8 53.78 + 19.7 60.03 + 21.7 48.85 + 8.7 54.57 + 10.6 LLM 52.81 + 17.5 60.17 + 21.9 52.45 + 16.7 55.77 + 13.0 52.45 + 16.7 59.45 + 20.5 LLMC 56.89 + 26.6 66.26 + 34.3 57.83 + 28.7 66.98 + 35.8 57.03 + 26.9 65.12 + 32.1 Children Origin 2.25 + 0.0 3.00 + 0.0 2.25 + 0.0 3.00 + 0.0 2.25 + 0.0 3.00 + 0.0 Num 15.87 + 605.3 16.01 + 433.7 15.93 + 608.0 16.68 + 456.0 15.75 + 600.0 15.80 + 426.7 NumC 19.32 + 758.7 18.87 + 529.0 19.07 + 747.6 18.10 + 503.3 18.13 + 705.8 17.50 + 483.3 LLM 17.94 + 697.3 17.82 + 494.0 18.82 + 736.4 19.98 + 566.0 18.06 + 702.7 18.31 + 510.3 LLMC 19.69 + 775.1 18.64 + 521.3 22.16 + 884.9 24.50 + 716.7 21.81 + 869.3 22.99 + 666.3
Table 2: Ablation Studies. Node classification under three variants of SaVe-TAG. Origin uses no augmentation; Num and LLM apply numeric and LLM-based interpolation, respectively. Subscript C indicates the use of confidence-based edge generation versus naive edge duplication otherwise.
5.4 Enhancement in Long-tailed Graph Learning (RQ3)
Figure 6: Node Classification Improvement.
@@ 5.4 Enhancement in Long-tailed Graph Learning (RQ3)
Table 3 shows that our method, particularly SaVe-TagS, outperforms prior baselines in node classification on long-tailed text-attributed graphs.
Moreover, our method effectively narrows the disparity between the head (majority) and tail (minority) classes. As shown in Figure 6a, across all benchmarks, SaVe-TagS achieves a smaller accuracy gap than numeric interpolation, underscoring its effectiveness in mitigating class imbalance in long-tailed learning.
@@
We further perform a sensitivity analysis to evaluate the stability of SaVe-Tag  under varying imbalance ratios. As shown in Figure 6b, SaVe-TagS exhibits smaller fluctuations compared to other baselines as imbalance ratios vary. Additional results are shown in the Appendix  A.
Method Cora Pubmed Citeseer Acc F1 GMean bAcc Acc F1 GMean bAcc Acc F1 GMean bAcc SMOTE 71.8$\pm$2.7 70.8$\pm$2.1 82.7$\pm$1.7 71.8$\pm$2.7 68.5$\pm$6.2 65.5$\pm$10.1 75.9$\pm$4.8 68.5$\pm$6.2 50.7$\pm$3.0 47.9$\pm$3.1 67.6$\pm$2.2 50.7$\pm$3.0 Oversampling 72.9$\pm$1.3 72.0$\pm$0.9 83.4$\pm$0.9 72.9$\pm$1.3 58.5$\pm$6.3 50.4$\pm$6.5 68.1$\pm$5.1 58.5$\pm$6.3 52.0$\pm$4.9 49.6$\pm$4.7 68.5$\pm$3.6 52.0$\pm$4.9 EmbedSMOTE 70.0$\pm$1.6 69.8$\pm$1.4 81.5$\pm$1.0 70.0$\pm$1.6 63.4$\pm$7.6 60.3$\pm$9.7 72.0$\pm$5.9 63.4$\pm$7.6 47.5$\pm$2.0 45.4$\pm$2.5 65.2$\pm$1.5 47.5$\pm$2.0 GraphSMOTET 74.4$\pm$4.0 74.2$\pm$4.3 84.4$\pm$2.6 74.4$\pm$4.0 67.9$\pm$5.0 67.8$\pm$5.0 75.5$\pm$3.9 67.9$\pm$5.0 50.5$\pm$2.8 48.1$\pm$3.2 67.4$\pm$2.1 50.5$\pm$2.8 GraphSMOTEO 74.3$\pm$3.7 74.1$\pm$4.9 84.3$\pm$3.0 74.3$\pm$4.7 69.7$\pm$5.0 69.9$\pm$4.8 76.9$\pm$3.9 69.7$\pm$5.0 52.0$\pm$1.6 49.8$\pm$1.8 68.6$\pm$1.2 52.0$\pm$1.6 MixupForGraph 66.4$\pm$2.7 63.32$\pm$2.7 80.1$\pm$1.6 68.0$\pm$2.5 69.6$\pm$0.8 68.7$\pm$1.1 75.9$\pm$1.1 68.7$\pm$1.6 63.2$\pm$1.9 59.4$\pm$1.9 75.1$\pm$1.8 60.9$\pm$2.5 HierTail  [^1] Modified code while running on PubMed dataset to prevent out-of-memory issues due to scalability. 72.8$\pm$5.3 73.3$\pm$4.7 83.3$\pm$3.4 72.8$\pm$5.3 72.4$\pm$8.2 72.4$\pm$7.7 79.0$\pm$6.3 72.5$\pm$8.0 49.6$\pm$9.6 44.5$\pm$11.3 66.5$\pm$7.4 49.6$\pm$9.6 LTE4G  [^1] Modified code while running on PubMed dataset to prevent out-of-memory issues due to scalability. 74.2$\pm$3.8 74.1$\pm$3.8 84.3$\pm$2.4 74.2$\pm$3.8 72.9$\pm$1.5 72.6$\pm$1.9 79.4$\pm$1.2 72.9$\pm$1.5 51.4$\pm$3.8 49.2$\pm$3.7 68.1$\pm$2.8 51.4$\pm$3.8 SaVe-TagO 72.1$\pm$0.1 70.9$\pm$0.1 84.5$\pm$0.3 74.8$\pm$0.4 72.8$\pm$0.3 73.0$\pm$0.4 80.2$\pm$0.2 75.0$\pm$0.2 69.9$\pm$0.2 66.8$\pm$0.2 79.9$\pm$0.1 67.8$\pm$0.2 SaVe-TagM 73.0$\pm$0.4 71.8$\pm$0.3 85.1$\pm$0.2 75.8$\pm$0.4 71.3$\pm$0.2 71.2$\pm$0.3 79.1$\pm$0.1 73.6$\pm$0.3 69.5$\pm$0.3 65.9$\pm$0.4 79.3$\pm$0.3 66.9$\pm$0.5 SaVe-TagS 76.4$\pm$0.3 74.1$\pm$0.3 85.8$\pm$0.2 76.6$\pm$0.3 76.3$\pm$0.7 76.4$\pm$0.8 82.5$\pm$0.5 77.6$\pm$0.7 69.9$\pm$0.3 66.9$\pm$0.3 80.1$\pm$0.3 68.1$\pm$0.4
Table 3: Comparison of SaVe-Tag with previous long-tail graph learning baselines.
5.5 Ablation Studies
@@ 6 Discussion
Here we outline relevant concepts and advances; see Appendix D for details.
  • Vicinal Risk Minimization (VRM): VRM [35] extends ERM [5] by estimating risk from local vicinity distributions, enhancing model generalization. It underlies methods like Mixup [12] and SMOTE [7], which create synthetic samples via interpolation.
  • Long-tailed Graph Learning: In graph learning, long-tailed node classification is typically addressed via hierarchical task grouping like HierTail [2] and expert models like LTE4G [3].
  • Large Language Model (LLM) Data Augmentation: Generally, LLMs augment data by generating contextually relevant text via zero-shot or few-shot instruction-based prompts [36, 10].
@@
While LLM-based interpolation offers clear benefits, it is more time- and resource-intensive than numeric embedding methods. Accordingly, we provide a cost analysis in Appendix C. Despite the overhead, we find the trade-off worthwhile: our framework supports small, open-source language models, requires no extensive training or tuning, and applies LLMs only once during augmentation—allowing reuse across tasks. Most importantly, our method consistently outperforms baselines, validating the effectiveness of LLM interpolation in practice.
@@ Appendix A Additional experimental results
A.1 Semantic-aware Benefits.
UMAP Visualization.
Figure 7 shows the UMAP visualization across all six datasets, with the boundary of the original data circled. As we can see, for all the datasets, numeric interpolated samples tend to stay within the original boundary, while LLM-generated samples often extend beyond it, enlarging the VRM decision region.
Figure 7: UMAP projection of the original and interpolated samples across various benchmarks—the original data boundary is outlined, with numeric samples mostly staying within and LLM-generated samples extending beyond.
Boundary Coverage Rate (BCR) and Boundary Proximity Score (BPS).
Figure 8: Boundary Coverage Rate (BCR) and Boundary Proximity Score (BPS) for numeric (blue) vs. LLM (yellow) interpolation in Interp+Orig and Interp-Only settings across six datasets; the red dashed line marks the original baseline.
Figure 8 presents the evaluation results using BCR and BPS metrics across all datasets. We report performance under three settings: original data, the combination of original and interpolated samples (Interp+Orig), and interpolated samples alone (Interp-Only).
Across all benchmarks, our LLM-based interpolation consistently outperforms numeric interpolation on both metrics, suggesting that LLM-generated samples align more effectively with decision boundaries. Particularly in some cases, LLM-based interpolation even exceeds the original data in the BCR, highlighting its potential to extend and enrich the original decision boundary.
Figure 9: Node Classification Improvement.
In-Class Rate (ICR).
We train a separate MLP classifier on each dataset using a balanced set of class samples to assess class consistency. The average In-Class Rate (ICR)—the proportion of augmented samples classified into their intended class—is reported in Figure 9.
Numerical interpolation achieves 100% ICR on Cora, PubMed, and Citeseer, and nearly 100% on others, reflecting high class consistency. In contrast, LLM-based samples show lower ICRs, indicating proximity to decision boundaries that causes occasional misclassification by MLP.
Node classification Performance
We evaluate the impact of LLM-based and numeric interpolation on node classification across multiple benchmarks. Results are presented in Figure 10. Embed utilizes LLaMA 3.2-1B embeddings for interpolation, while SimCSE and SBERT serve as embedding-based numeric baselines. Across all datasets, the our LLM-based approach with LLaMA 3.2-1B consistently outperforms all numeric interpolation baselines.
Figure 10: Node classification performance (F1) under various implementation of SaVe-TagS. Embed uses Llama3.2-1B embeddings for interpolation. SimCSE and SBERT are compared as alternative embedding baselines. Our LLM-based method with Llama3.2-1B consistently outperforms.
A.2 Necessity of Topology-aware Filters
Figure 11: Impact of average assigned edges on performance for Photo (a), Children (b), and Computer (c). We report changes from the baseline ($n{=}1$) across four metrics: $\Delta$Acc, $\Delta$F1, $\Delta$bAcc, and $\Delta$GMean.
We instantiate Definition 4.7 using a top–$k$ rule, where each synthetic node retains its $k$ highest-confidence edges. The value of $k$ is set to #$\#augmented\_nodes\times n$, and we vary $n$ to study its effect on node isolation and classification performance. Figure 11 reveals a bell-shaped performance curve with respect to $\log(n)$ on Photo, Computer, and Children, measured relative to $n{=}1$. More discussion can be found in Section 5.3.
A.3 Enhancement in Long-tailed Graph Learning.
Stability.
We also adjusted the imbalance ratio, selecting from $\{0.1,0.2,0.3,0.4,0.5\}$, to examine how our model performs under different levels of imbalance. As shown in Figure 12, SaVe-TagS (the bold line) exhibits less performance drop as the imbalance ratio decreases, demonstrating less fluctuation compared to other baselines.
Figure 12: F1 score (%) versus imbalance ratio on (a) Cora, (b) PubMed, and (c) Citeseer, comparing other baselines of long-tailed graph learning and with our SaVe-TagS (the bold line).
@@ A.2 Necessity of Topology-aware Filters
To assess the generalization of our method across different classifier architectures, we conduct additional ablation studies using one GNN-based backbone (SAGE) and one non-graph-based model (MLP). Since MLP does not incorporate structural information (i.e., it lacks neighborhood aggregation), the confidence-based edge assignment has no effect. Therefore, we report only the results without the subscript C. As shown in Table 4, SaVe-Tag  effectively addresses long-tailed classification with both GNNs and non-graph-based models. Its strong performance with MLP highlights its potential for text classification and broader applications.
@@
Model Method Cora PubMed Citeseer Photo Computer Children F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc SaVe-TagS MLP Origin 50.31 52.34 25.51 42.59 61.70 64.64 4.25 7.48 3.09 4.61 1.06 2.03 Num 58.28 61.35 64.63 68.62 61.70 64.64 34.96 34.74 28.19 31.77 15.31 14.85 LLM 64.13 66.80 72.77 72.07 64.04 66.25 43.26 40.79 37.65 40.26 16.64 16.83 SAGE Origin 67.5 70.2 60.7 62.4 52.9 55.5 47.4 44.2 36.5 38.0 2.5 4.0 Num 70.6 73.1 68.3 68.6 65.0 68.0 54.2 52.1 43.1 47.3 18.7 18.6 NumC 70.3 72.4 69.2 69.6 65.8 69.4 50.6 48.9 49.3 54.7 19.4 18.7 LLM 72.5 74.5 72.5 72.5 67.4 70.0 57.7 56.0 48.6 53.7 20.1 20.8 LLMC 74.0 76.4 76.4 76.3 66.6 69.3 57.6 56.1 52.0 59.1 20.3 20.3 SaVe-TagO MLP Origin 50.31 52.34 25.51 42.59 61.70 64.64 4.25 7.48 3.09 4.61 1.06 2.03 Num 57.85 60.97 66.06 66.46 60.23 63.38 35.47 35.01 28.65 32.32 15.89 15.32 LLM 63.02 64.12 70.25 69.88 64.68 67.07 43.34 42.15 38.18 42.74 18.65 20.38 SAGE Origin 67.5 70.2 60.7 62.4 52.9 55.5 47.4 44.2 36.5 38.0 2.5 4.0 Num 70.6 73.8 69.0 69.3 65.3 68.6 54.6 53.5 43.0 47.1 18.5 19.0 NumC 69.1 70.9 68.3 68.6 65.5 68.9 49.0 46.1 42.9 46.0 18.6 19.1 LLM 72.1 73.7 71.4 71.6 67.5 70.3 57.8 56.1 48.1 53.9 20.9 21.8 LLMC 69.2 70.1 72.5 72.3 66.2 69.0 58.0 57.3 52.5 59.9 22.5 23.8 SaVe-TagM MLP Origin 50.31 52.34 25.51 42.59 61.70 64.64 4.25 7.48 3.09 4.61 1.06 2.03 Num 61.66 62.89 67.88 67.77 61.06 63.57 36.98 35.39 29.25 33.24 15.95 15.95 LLM 62.94 64.47 71.12 70.56 64.52 66.64 44.84 42.28 38.47 41.47 18.80 20.09 SAGE Origin 67.5 70.2 60.7 62.4 52.9 55.5 47.4 44.2 36.5 38.0 2.5 4.0 Num 71.9 74.0 69.8 70.0 64.8 68.7 56.0 54.1 46.0 51.4 19.4 19.8 NumC 69.8 70.1 69.9 69.7 64.3 68.6 53.6 51.7 46.6 51.4 19.2 19.6 LLM 72.3 74.2 71.5 71.3 68.0 70.8 58.4 56.7 49.1 54.8 21.7 23.1 LLMC 70.5 72.5 75.7 75.2 67.8 70.9 58.6 57.8 50.6 57.4 23.9 26.5
@@
Table 4: Ablation Studies. Node classification performance using SAGE and MLP under three variants of SaVe-Tag. Origin denotes no augmentation; Num and LLM apply numeric and LLM-based interpolation, respectively. The subscript C indicates the use of confidence-based edge generation, while its absence denotes naive edge duplication. The best results for both MLP and SAGE are emboldened, and the second-best results for SAGE are underlined.

### v5

@@ 4.3. Necessity of Confidence-based Edge Assignment (RQ2)
We instantiate Definition 3.4 using a top–$k$ rule, where each synthetic node retains its $k$ highest-confidence edges. The value of $k$ is set to $augmented\_nodes\times n$, and we vary $n$ to study its effect on node isolation and classification performance. The left part of Figure 7 shows that the number of isolated nodes ($\log$-scaled) drop as $\log(n)$ increases. The right part of Figure 7 reveals bell-shaped performance curves with respect to $\log(n)$ on the Photo, Computer, and Children datasets, measured relative to the baselines of n=1n{=}1. These trends support our hypothesis:
  • Too small $k$ $\rightarrow$ isolation. Many synthetic nodes become singletons (Theorem 3.9), so the minority manifold is under-populated and performance suffers.
  • Too large $k$ $\rightarrow$ over-smoothing. Excessive connectivity causes repeated averaging in Eq.(2), pushing node representations toward a global mean and resulting in over-smoothing(Li et al., 2024a).
@@
We compare two edge assignment strategies for augmented nodes: edge duplication from the target node (without the ‘C{C}’ subscript) and edge assignment with our confidence-based function (denoted by ‘C{C}’). Table 3 shows that our confidence-based edge assignment consistently outperforms naive edge duplication under LLM-based interpolation. A similar trend holds for numeric interpolation, with the exception of Cora and Photo, likely due to the generation of off-manifold samples that our confidence mechanism successfully isolates.
Table 2. Comparison of SaVe-TAG with previous long-tail graph learning baselines.
[…]
  • 11footnotemark: 1
    Modified code while running on PubMed dataset to prevent out-of-memory issues due to scalability.
4.4. Enhancement in Long-tailed Graph Learning (RQ3)
Figure 8. Node Classification Enhancement. Left: SaVe-TAGS achieves a smaller Head – Tail accuracy gap than numeric interpolation, indicating better class balance. Right: SaVe-TAGS (bold) maintains more stable performance with less fluctuation across imbalance ratios.
Table 3. Ablation Studies. Node classification under three variants of SaVe-TAG. Origin uses no augmentation; Num and LLM apply numeric and LLM-based interpolation, respectively. Subscript C indicates the use of confidence-based edge generation versus naive edge duplication otherwise.
Dataset Method SaVe-TAGS SaVe-TAGM SaVe-TAGO F1 ACC F1 ACC F1 ACC Average across all 6 datasets Origin 45.00 47.55 45.00 47.55 45.00 47.55 Num 54.07 55.78 55.09 56.64 54.32 56.09 NumC 54.86 56.57 56.37 57.54 54.35 55.64 LLM 56.79 58.49 56.90 58.18 56.39 58.32 LLMC 58.85 61.03 59.02 61.74 59.24 62.05
@@ 4.4. Enhancement in Long-tailed Graph Learning (RQ3)
Table 2 shows that our method, particularly SaVe-TAGS, outperforms prior baselines in node classification on long-tailed text-attributed graphs. Moreover, our method effectively narrows the disparity between the head (majority) and tail (minority) classes. As shown in the left part of Figure 8, across all benchmarks, SaVe-TAGS consistently achieves a smaller disparity (the average accuracy gap between head and tail classes) than numeric interpolation, underscoring its effectiveness in mitigating class imbalance in long-tailed learning.
@@
We further perform a sensitivity analysis to evaluate the stability of SaVe-TAG  under varying imbalance ratios. As shown in the right part of Figure 8, SaVe-TAGS exhibits smaller fluctuations compared to other baselines as imbalance ratios vary.
4.5. Ablation Studies
@@ 5. Discussion
Here we outline relevant research topics. Vicinal Risk Minimization (VRM): VRM (Chapelle et al., 2000a) extends ERM (Vapnik and Vapnik, 1998) by estimating risk from local vicinity distributions, enhancing model generalization. It underlies methods like Mixup (Zhang et al., 2018) and SMOTE (Chawla et al., 2002), which create synthetic samples via interpolation. Long-tailed Graph Learning: In graph learning, long-tailed node classification is typically addressed via hierarchical task grouping like HierTail (Wang et al., 2024) and expert models like LTE4G (Yun et al., 2022). Large Language Model (LLM) Data Augmentation: Generally, LLMs augment data by generating contextually relevant text via zero-shot or few-shot instruction-based prompts (Li et al., 2024b, 2023).
@@
While LLM-based interpolation offers clear benefits, it is more resource-intensive than numeric embedding methods. We include a cost analysis in Appendix C. Despite the overhead, the trade-off is worthwhile: our framework supports small open-source language models, needs no extensive training/tuning, and applies LLMs only once during augmentation—allowing reuse across tasks. Most importantly, it consistently outperforms baselines, confirming the practical value of LLM interpolation.
@@ 6. Conclusion
Acknowledgements
This research is supported by the National Science Foundation (NSF) under grant numbers IIS2239881 and ECCS2325417.
@@ Appendix A Additional Experimental Results
To assess the generalization of our method across different classifier architectures, we conduct additional ablation studies using one GNN-based backbone (SAGE) and one non-graph-based model (MLP). Since MLP does not incorporate structural information (i.e., it lacks neighborhood aggregation), the confidence-based edge assignment has no effect. Therefore, we report only the results without the subscript C. Table 4 shows SaVe-TAG  effectively addresses long-tailed classification with both GNNs and MLP; its strong performance with MLP highlights its potential for text classification and broader applications.
Table 4. Ablation Studies. Node classification performance using SAGE and MLP under three variants of SaVe-TAG.
@@
Model Method Cora PubMed Citeseer Photo Computer Children
F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc
SaVe-TAGS MLP Origin 50.31 52.34 25.51 42.59 61.70 64.64 4.25 7.48 3.09 4.61 1.06 2.03
Num 58.28 61.35 64.63 68.62 61.70 64.64 34.96 34.74 28.19 31.77 15.31 14.85
LLM 64.13 66.80 72.77 72.07 64.04 66.25 43.26 40.79 37.65 40.26 16.64 16.83
SAGE Origin 67.5 70.2 60.7 62.4 52.9 55.5 47.4 44.2 36.5 38.0 2.5 4.0
Num 70.6 73.1 68.3 68.6 65.0 68.0 54.2 52.1 43.1 47.3 18.7 18.6
NumC 70.3 72.4 69.2 69.6 65.8 69.4 50.6 48.9 49.3 54.7 19.4 18.7
LLM 72.5 74.5 72.5 72.5 67.4 70.0 57.7 56.0 48.6 53.7 20.1 20.8
LLMC 74.0 76.4 76.4 76.3 66.6 69.3 57.6 56.1 52.0 59.1 20.3 20.3
SaVe-TAGO MLP Origin 50.31 52.34 25.51 42.59 61.70 64.64 4.25 7.48 3.09 4.61 1.06 2.03
Num 57.85 60.97 66.06 66.46 60.23 63.38 35.47 35.01 28.65 32.32 15.89 15.32
LLM 63.02 64.12 70.25 69.88 64.68 67.07 43.34 42.15 38.18 42.74 18.65 20.38
SAGE Origin 67.5 70.2 60.7 62.4 52.9 55.5 47.4 44.2 36.5 38.0 2.5 4.0
Num 70.6 73.8 69.0 69.3 65.3 68.6 54.6 53.5 43.0 47.1 18.5 19.0
NumC 69.1 70.9 68.3 68.6 65.5 68.9 49.0 46.1 42.9 46.0 18.6 19.1
LLM 72.1 73.7 71.4 71.6 67.5 70.3 57.8 56.1 48.1 53.9 20.9 21.8
LLMC 69.2 70.1 72.5 72.3 66.2 69.0 58.0 57.3 52.5 59.9 22.5 23.8
SaVe-TAGM MLP Origin 50.31 52.34 25.51 42.59 61.70 64.64 4.25 7.48 3.09 4.61 1.06 2.03
Num 61.66 62.89 67.88 67.77 61.06 63.57 36.98 35.39 29.25 33.24 15.95 15.95
LLM 62.94 64.47 71.12 70.56 64.52 66.64 44.84 42.28 38.47 41.47 18.80 20.09
SAGE Origin 67.5 70.2 60.7 62.4 52.9 55.5 47.4 44.2 36.5 38.0 2.5 4.0
Num 71.9 74.0 69.8 70.0 64.8 68.7 56.0 54.1 46.0 51.4 19.4 19.8
NumC 69.8 70.1 69.9 69.7 64.3 68.6 53.6 51.7 46.6 51.4 19.2 19.6
LLM 72.3 74.2 71.5 71.3 68.0 70.8 58.4 56.7 49.1 54.8 21.7 23.1
LLMC 70.5 72.5 75.7 75.2 67.8 70.9 58.6 57.8 50.6 57.4 23.9 26.5
@@
  • Origin denotes no augmentation; Num and LLM apply numeric and LLM-based interpolation, respectively. The subscript C indicates confidence-based edge generation and its absence denotes naive edge duplication. The best results for both MLP/SAGE are in bold, and the second-best results for SAGE are underlined.
Table 5. Performance with additional LLM variants.
Model Cora PubMed Citeseer
Acc F1 GMean bAcc Acc F1 GMean bAcc Acc F1 GMean bAcc
Qwen2.5‑7B‑Instruct 73.6 71.9 85.3 76.1 74.2 74.2 81.0 75.7 72.1 67.9 80.4 68.4
Mistral‑8B‑Instruct‑2040 74.3 72.9 85.2 75.8 73.5 73.5 80.8 75.7 71.1 66.7 79.7 67.3
LLM Variants.
To further examine model variation, we conducted experiments with Qwen2.5‑7B‑Instruct and Mistral‑8B‑Instruct‑2040. The results using SaVe‑TAG are shown in Table 5.